You are considering a stock investment in one of two firms (NoEquity, Inc. and NoDebt, Inc.), both of which operate in the same industry and have identical operating income of $20.5 million. NoEquity, Inc. finances its $70 million in assets with $69 million in debt (on which it pays 10 percent interest annually) and $1 million in equity. NoDebt, Inc. finances its $70 million in assets with no debt and $70 million in equity. Both firms pay a tax rate of 30 percent on their taxable income. Calculate the net income and return on assets for the two firms. (Enter your dollar answers in millions of dollars. Round all answers to 2 decimal places.) Answer is complete but not entirely correct NoEquity 3.40million 14.35 million 48.50 % NoDebt Net income Return on assets 20.50 %
Answer #1

No equity:

Interest = 69 * 0.1 = 6.9 million

Net income = (EBIT - Interest)(1 - tax)

Net income = (20.5 - 6.9)(1 - 0.3)

Net income = 9.52 million

ROE = (Net income / assets) * 100

ROE = (9.52 / 70) * 100

ROE = 13.6%

No debt:

Net income = (EBIT)(1 - tax)

Net income = (20.5)(1 - 0.3)

Net income = 14.35 million

ROE = (Net income / assets) * 100

ROE = (14.35 / 70) * 100

ROE = 20.50%